I really need help on this problem. Its very confusing to me.
The equation that expressed Celsius temperature C in terms of Fehrenheit temperature F is linear. Water freezes at 32 degreesF or 0 degreesC and boils at 212 degreesF or 100 degreesC. Find the linear equation of C in terms of F. Please show me a complete solution so I know how to do this for future reference. Thank You!!!!

Answer by stanbon(57290)

You want C in terms of F or C(F)
You have two points: (32,0) and (212,100)
The slope is (100-0)/(212-32)= 100/180 = 5/9
Your equation looks line C=mF+b
You have values for C,m, and F; solve for b.
Using (32,0) you get 0=(5/9)(32)+b
b=-17.78
So, C=(5/9)F-17.78
It is easier to remember the following which converts
from Celsius to Fahrenheit: F=(9/5)C+32
Cheers,
Stan H.